OpenStudy (mayaal):

Could I find the two missing angle measures if I know some of the side lengths of a right triangle?

OpenStudy (mayaal):

I NEED AN EXAMPLE TOO

OpenStudy (campbell_st):

yes, but you will need 2 side lengths... find 1 angle and then use angle sum of a triangle to find the other

OpenStudy (campbell_st):

 find B or A using the tan ratio then find the other by angle sum of a triangle

OpenStudy (anonymous):

If it's a right triangle, all that is necessary is two side lengths, because you can solve for the remaining side. Then, simply use the law of cosine to find the angles.

OpenStudy (mayaal):

what do u mean by the angle sum of a triangle?

OpenStudy (mayaal):

so A=1.667DEGREES AND B=0.6 DEGREES?

OpenStudy (campbell_st):

well the angles in a triangle add to 180 degrees if 1 angle is know, 90 and you find another using trig. you can simply use the angle sum to find the 3rd angle, rather than use trig again.

OpenStudy (anonymous):

In his example, the remaining side would be \[\sqrt{34}\]. Then, to find angle A: \[25=9+34-2*3*\sqrt{34}*\cos(A)\] Use basic algebra to find A.

OpenStudy (mayaal):

oh,i get it thanku so much,both of u!!!

OpenStudy (campbell_st):

why use the law of consine in a right triangle when you can use the tan ratio tan(B) = 3/5

OpenStudy (anonymous):

Because the law of cosine is super cool.

OpenStudy (campbell_st):

lol... ok....

OpenStudy (mayaal):

hahaha

OpenStudy (campbell_st):

but more calculations leads to more errors

OpenStudy (anonymous):

Maybe for YOU, Smart Cookie. :P